OpenStudy (anonymous):

Simplify: 3+2i over 4-i

Nnesha (nnesha):

multiply top and bottom of the fraction by conjugate of the denominator

Nnesha (nnesha):

what is the `conjugate` of 4-i ? do you know ?

OpenStudy (anonymous):

Welcome to openstudy!

Nnesha (nnesha):

thanks :=)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

I don't know the conjugate ;~;

Nnesha (nnesha):

here is an example the conjugate of a+bi is `a-bi` so change the sign of the imaginary term

OpenStudy (anonymous):

I don't get how to do that either...

Nnesha (nnesha):

conjugate of a+bi is `a-bi` right so what's the difference btw `a+bi` and `a-bi` ?

OpenStudy (anonymous):

The + and -

Nnesha (nnesha):

yes right so you should change the sign of imaginary number `i` = imaginary

Nnesha (nnesha):

so what's the conjugate of 4-i ?

OpenStudy (anonymous):

4+i?

Nnesha (nnesha):

yes right! now multiply \[\huge\rm \frac{ 3+2i }{ 4-i} \times \frac{ 4+i }{ 4+i }\]

OpenStudy (anonymous):

12+3i over 16?

Nnesha (nnesha):

show ur steps please how did you get that ?

OpenStudy (anonymous):

I don't know how I did that. I have no idea how i did that.

Nnesha (nnesha):

then why did you post that ?

OpenStudy (anonymous):

I was hoping it was right..

Nnesha (nnesha):

familiar with the foil method ?

OpenStudy (anonymous):

Yeah.

Nnesha (nnesha):

i know that's one of ur option.

Nnesha (nnesha):

alright good apply foil method \[\huge\rm \frac{ (3+2i)(4+i) }{ (4-i)(4+i) }\] (3+2i)(4+i) = ??

OpenStudy (anonymous):

I think it's 12+14i? That's what I got..

Nnesha (nnesha):

okay plz show ur work so i can look over to find out mistakes :=)

Nnesha (nnesha):

that would be great!

OpenStudy (anonymous):

(3+2i)(4+i) 12+3i+8i+3i 12+6i+8i 12+14i

Nnesha (nnesha):

nice thanks so 2i times i = ??

OpenStudy (anonymous):

Not 3i?

OpenStudy (anonymous):

OH OH -2

Nnesha (nnesha):

yes right i times i= i^2 and i^2 =-1 so 2(-1) = -1 now fix it

Nnesha (nnesha):

so it's \[12+3i+8i-2\]

OpenStudy (anonymous):

So I was mostly right. Now what.

Nnesha (nnesha):

now combine like terms 12+3i+8i-2 = ?

OpenStudy (anonymous):

10+11i

Nnesha (nnesha):

looks right! now simplify the denominator (4-i)(4+i)= ?

OpenStudy (anonymous):

17!

Nnesha (nnesha):

right that's denominator so what would be the final answer ?

OpenStudy (anonymous):

\[\frac{ 10+11i}{ 17}\]

Nnesha (nnesha):

looks good ! great job !

OpenStudy (anonymous):

Thank you~

Nnesha (nnesha):

np :=)